OpenStudy (anonymous):

Sales tax charged by the state of Virginia is 4.5%. This is 2.5% less than the sales tax charged by the state of Rhode Island. What does Rhode Island charge for sales tax? Write an equation to model each situation. Then Solve.

OpenStudy (anonymous):

Please I need help!

OpenStudy (phi):

This is an ambiguous question. Normally 2.5% less than x would mean x-0.025*x. For example 2% less than 100 would be 100-0.02*100= 100-2=98. However, here they actually mean Rhode Island's sales tax is 4.5%+2.5%=7%.

OpenStudy (anonymous):

Thanks a whole bunch. I was playing school and was teach first step equations. I had a equation about it